OpenStudy (anonymous):

Given the equation v=-3000t+21000 can be used to find the value of V of a car at the end of t years, is it true that the car will be worth 200 after 7 years 200 dollars that is

OpenStudy (anonymous):

v = -3000t + 21000 (200) = -3000t + 21000 200 - 21000 = -3000t -19000 = -3000t t = 6.33333... No, it is not true because the car will be worth $200 in 6.33333... years, not 7 years. Alternatively, you could plug in t = 7 to find out what the value will be in 7 years. v = -3000(7) + 21000 v = -21000 + 21000 v = 0 Lol I don't think the car would ever be worth exactly $0, but for purposes of this question it is. So therefore, no, the car wouldn't be worth $200 in 7 years because it would be worth $0 in 7 years, not $200.

OpenStudy (anonymous):

LOL! YES funny that a car is worth 0! But thank you so much!